# FastCuRL: Curriculum Reinforcement Learning with Stage-wise Context Scaling for Efficient Training R1-like Reasoning Models

### v5

@@ Abstract
Improving training efficiency continues to be one of the primary challenges in large-scale Reinforcement Learning (RL). In this paper, we investigate how context length and the complexity of training data influence the RL scaling training process of R1-distilled reasoning models, e.g., DeepSeek-R1-Distill-Qwen-1.5B. Our experimental results reveal that: (1) simply controlling the context length and curating the training data based on the input prompt length can effectively improve the training efficiency of RL scaling, achieving better performance with more concise CoT; (2) properly scaling the context length helps mitigate entropy collapse; and (3) carefully choosing the context length facilitates achieving efficient LLM training and reasoning. Inspired by these insights, we propose FastCuRL, a curriculum RL framework with stage-wise context scaling to achieve efficient LLM training and reasoning. Extensive experimental results demonstrate that FastCuRL-1.5B-V3 significantly outperforms state-of-the-art reasoning models on five competition-level benchmarks and achieves 49.6% accuracy on AIME 2024. Furthermore, FastCuRL-1.5B-Preview surpasses DeepScaleR-1.5B-Preview on five benchmarks while only using a single node with 8 GPUs and a total of 50% of training steps. The code, training data, and models have been publicly released [^1] https://github.com/nick7nlp/FastCuRL.
@@ 1 Introduction
A primary finding from recent breakthroughs, exemplified by DeepSeek-R1 DeepSeek-AI (2025), reveals a scaling phenomenon in the training process of Reinforcement Learning (RL). Inspired by these findings, training LLMs through scaling RL has recently emerged as a promising paradigm for addressing complex reasoning tasks and many valuable research endeavours Face (2025); Hu et al. (2025); Zeng et al. (2025); Liu et al. (2025) have emerged to explore and replicate reasoning models akin to DeepSeek-R1 (for example, starting from R1-distilled or pre-trained models) by extending the generation length of CoT.
However, excessively long CoT responses significantly increase computational overhead during model training and deployment. Moreover, recent studies Yeo et al. (2025); Wu et al. (2025); Team (2025a) have identified an inherent overthinking phenomenon in reasoning models, which includes irrelevant details and repetitive thinking patterns. This kind of information leads to inefficient use of computational resources and undermines reasoning accuracy, which causes models to stray from valid logical pathways, resulting in incorrect answers. To this end, recent studies Team (2025a); Luo et al. (2025); Liu et al. (2025); Yu et al. (2025) focus on efficient reasoning for optimizing the model to generate more concise solutions. Among them, DeepScaleR Luo et al. (2025) propose to iteratively increase the context length from 8K to 24K to train the DeepSeek-R1-Distill-Qwen-1.5B model toward more concise reasoning, outperforming OpenAI’s o1-preview OpenAI (2024). By observing the training logs [^2] https://github.com/agentica-project/rllm of DeepScaleR in Figure 2, we find two issues:
@@
To this end, in this paper, we investigate how the model’s context length and the complexity of the training dataset influence the training process of R1-like reasoning models. Motivated by our observations, we propose FastCuRL, a simple yet efficient Curriculum Reinforcement Learning framework with a stage-wise context scaling strategy to improve the RL training efficiency and achieve concise CoT reasoning for R1-like reasoning models. Experimental results demonstrate that our model FastCuRL-1.5B-V3 outperforms recent state-of-the-art reasoning baselines across five competition-level benchmarks, AIME 2024, AMC 2023, MATH 500, Minerva Math, and OlympiadBench. Furthermore, our model FastCuRL-1.5B-Preview surpasses DeepScaleR-1.5B-Preview on five benchmarks and only uses 50% training steps on a single node with 8 GPUs. We hope the findings presented in this paper, the models we have released, and the open-sourced code will benefit future research.
@@ 2 Methodology
In this section, we introduce our method in detail, which consists of three main parts: (1) a resource-efficient RL algorithm, (2) a complexity-aware, mathematics-focused dataset, and (3) a stage-wise context scaling approach. These three components aim to achieve efficient LLM reasoning.
@@ 2.1 Reinforcement Learning Algorithm
where the advantage $\mathrm{A}_{i}$ is computed from a group of rewards $\{r_{1},r_{2},...,r_{\mathrm{G}}\}$:
@@
2.2 Complexity-Aware Data Curation
@@ 3 Experiments
To investigate the research question described in Section 1—namely, how the model’s context length and the complexity of the training data influence the RL training process of R1-like reasoning models—we designed a set of experiments under computational resource constraints. We aim to analyze the training behavior of small LLMs and find practical insights. These experiments are intended not only to provide empirical evidence of performance gains but also to offer clear and actionable guidance for both future academic research and practical industry implementations.
@@ 3.1 Experimental Setup
3.2 Baselines
3.3 Benchmarks
3.4 Evaluation Metric
$\textsc{Pass@1}=\frac{1}{k}\sum_{i=1}^{k}p_{i},$ (4)
where $p_{i}$ is the correctness of the $i$-th response.
3.5 Main Processes and Results
In this section, we first validate the effectiveness of the complexity-aware data curation strategy. Then, we design a series of progressive experiments with varying context lengths and data complexities and analyze the experimental results.
3.5.1 Dataset Complexity Verification
To validate the effectiveness of complexity-aware data curation, we train three models with the same setting on L1, L2, and L3 under the 8K context length as seen from Figure 5, whether the experiment results meet expectations in clipping ratio, response length, and reward scores. These experimental results support our hypothesis that the more complex the problem, the longer the output the model needs to produce to arrive at a solution.
@@ 3.5.1 Dataset Complexity Verification
Table 3: Pass@1 performance across competition-level mathematics benchmarks.
@@
3.5.2 Multi-Stage Experimental Results
We conduct three sets of multi-stage experiments, with specific parameter settings shown in Table 3. These experiments include ones with 3, 4, and 5 training stages, respectively. The experimental results are presented in Table 2.
For the first set of experiments, Exp-3 achieves better performance compared to Exp-1, but it required more training steps (Exp-3 is trained based on the $\mathbf{L}$2 dataset twice). Therefore, by comparing the differences in effectiveness and the computational cost in terms of training steps, we select Exp-1 as the output of the first stage and adopt it as the base model for the second stage.
In the first set of experiments, we observe that the average response length in its final stage is between 6,000 and 7,000 tokens. Therefore, we test context lengths that are longer, shorter, and equal to the 24K context length. As shown in Table 2, setting the context length to 16K yielded the best performance, rather than longer contexts of 24K or 32K tokens. Therefore, we select Exp-6 as the base model for the third stage.
Inspired by the second set of experiments, we conduct a third set in which we set the context lengths to 24K, 16K, and 8 K. As shown in Table 2, the 16K context still achieves the best performance, but there is virtually no difference compared to the fourth stage. Analyzing this phenomenon, we find that during progressive context extension training, the model’s output length is initially constrained by the short context in the first stage. This constraint compresses the length of the thoughts but improves their quality. As the context increases in the second and third stages, the model begins to explore problems that require longer thought. However, this extension also introduces repetitive thought patterns. These repetitive patterns do not enhance the model’s reasoning capabilities; on the contrary, they may decrease the model’s exploratory efficiency, especially when the context length becomes excessively long. Therefore, further compressing the context length (as in the fourth stage) is necessary to improve the quality of the chain-of-thought and enhance the model’s exploratory efficiency.
In the third set of experiments, we observe that neither increasing nor decreasing the context length from 16K yields comparable effectiveness, suggesting that maintaining the context length at 16K produces optimal results. This phenomenon raises an important question: Does there exist an optimal "sweet spot" for context length in R1-like models? Specifically, for the DeepSeek-R1-Distill-Qwen-1.5B model, is 16K the precise optimal context length, or does it merely approximate the optimal value more closely than the alternative configurations of 8K and 24K?
To investigate this question, we conducted a comprehensive series of experiments examining the model’s behavior across various context lengths. We trained the model with different context length configurations while maintaining the entropy coefficient at $1\times 10^{-6}$ to observe entropy dynamics. As illustrated in Figure 6, our findings reveal a notable dichotomy in model behavior. For context lengths of 4K, 8K, and 12K, the entropy exhibits a rapid decline, suggesting a significant degradation in the model’s exploratory capabilities. In contrast, when the context lengths are configured at 16K, 20K, and 24K, the entropy demonstrates remarkable stability, converging to and maintaining a consistent value.
@@ 3.5.2 Multi-Stage Experimental Results
Table 5: Statistics of the responses of DeepSeek-R1-Distill-Qwen-1.5B and FastCuRL-1.5B-Preview.
@@
3.5.3 Overall Comparison Results
Table 3 present the overall Pass@1 performance of Qwen2.5-Math-7B-Instruct, DeepSeek-R1-Distill-Qwen-1.5B, Still-1.5B, Qwen2.5-7B-SimpleRL, rStar-Math-7B, Eurus-2-7B-PRIME, and DeepScaleR-1.5B-Preview. Specifically, our models achieve the best overall performance on five competition-level benchmarks.
Meanwhile, FastCuRL-1.5B-Preview has better generalization on the AMC 2023 and Minerva Math test sets than the baseline DeepScaleR-1.5B-Preview. Furthermore, as shown in Table 4, compared with the baseline DeepScaleR-1.5B-Preview, we only use 50% of the training steps during training and only one node with 8 GPUs, saving more than half of the training resources. Moreover, our model can achieve better results when using the same or more training steps.
3.5.4 Analyzing Generated Responses
@@
Figure 8: Entropy values of FastCuRL-1.5B-Preview across four RL training stages. Here, data points are smoothed using a running average (window size = 10) and plotted at uniform intervals.
Figure 7 compares DeepSeek-R1-Distill-Qwen-1.5B and FastCuRL-1.5B-Preview on the AIME 2024, measuring the average response length between correct and incorrect answers at the problem level to observe and analyse whether the long incorrect response is related to the difficulty of the problem. Across both models, incorrect answers (red bars) almost universally have greater average response lengths than correct answers (green bars). This suggests that models tend to generate more verbose content when producing incorrect answers, potentially reflecting "over-explanation" or "verbose reasoning" when the model is uncertain.
Figure 9: Performance comparison of training with and without Entropy loss at 8k context length.
@@ 4 Related Work
Advancements in RL methodologies have considerably enhanced the reasoning capabilities of LLMs. A pivotal development in this domain is OpenAI’s o1 OpenAI (2024), which employs RL training to promote the development of long CoT reasoning in LLMs. This approach has significantly enhanced performance on complex mathematical and programming benchmarks. Building upon this foundation, DeepSeek-R1 DeepSeek-AI (2025) demonstrates that pure RL post-training via Group Reinforcement Policy Optimization (GRPO), without needing supervised pre-training, can directly perform robust CoT reasoning capabilities. Notably, this method not only achieves performance competitive with o1 but also exhibits emergent behaviors such as self-verification and multi-step planning. Building on these advancements, the research community has been collectively working to study and apply DeepSeek-R1‘s methodology to enhance the reasoning capabilities of various sizes of language models, yielding remarkable progress, such as Face (2025); Luo et al. (2025); Zeng et al. (2025); Liu et al. (2025); Yu et al. (2025).

### v6

@@ Abstract
Improving training efficiency continues to be one of the primary challenges in large-scale Reinforcement Learning (RL). In this paper, we investigate how context length and the complexity of training data influence the RL scaling training process of R1-distilled reasoning models, e.g., DeepSeek-R1-Distill-Qwen-1.5B. Our experimental results reveal that: (1) simply controlling the context length and selecting the training data based on the input prompt length can effectively improve the training efficiency of RL scaling, achieving better performance with more concise CoT; (2) properly scaling the context length helps mitigate entropy collapse; and (3) carefully choosing the context length facilitates achieving efficient LLM training and reasoning. Inspired by these insights, we propose FastCuRL, a curriculum RL framework with stage-wise context scaling to achieve efficient LLM training and reasoning. Extensive experimental results demonstrate that FastCuRL-1.5B-V3 significantly outperforms state-of-the-art reasoning models on five competition-level benchmarks and achieves 49.6% accuracy on AIME 2024. Furthermore, FastCuRL-1.5B-Preview surpasses DeepScaleR-1.5B-Preview on five benchmarks while only using a single node with 8 GPUs and a total of 50% of training steps, as shown in Figure 1. The code, training data, and models have been publicly released [^1] https://github.com/nick7nlp/FastCuRL.
@@ 1 Introduction
A primary finding from recent breakthroughs, exemplified by DeepSeek-R1 DeepSeek-AI (2025), have revealed a notable scaling phenomenon in the training process of Reinforcement Learning (RL). Inspired by these findings, scaling RL for training LLMs has emerged as a promising paradigm for tackling complex reasoning tasks. This has sparked numerous valuable research efforts aimed at exploring and replicating reasoning models similar to DeepSeek-R1 Face (2025); Hu et al. (2025); Zeng et al. (2025); Liu et al. (2025). These initiatives typically begin with R1-distilled or pre-trained models and extend the generation length of CoT reasoning to enhance performance.
However, excessively long CoT responses significantly increase computational overhead during model training and deployment. Moreover, recent studies Yeo et al. (2025); Wu et al. (2025); Team (2025a) have identified an inherent overthinking phenomenon in reasoning models, which includes irrelevant details and repetitive thinking patterns. This kind of information leads to inefficient use of computational resources and undermines reasoning accuracy, which causes models to stray from valid logical pathways, resulting in incorrect answers. To this end, recent studies Team (2025a); Luo et al. (2025); Liu et al. (2025); Song and Zheng (2025) focus on efficient reasoning for optimizing the model to generate more concise CoT outputs. Among them, DeepScaleR Luo et al. (2025) propose to iteratively increase the context length from 8K to 24K to train the DeepSeek-R1-Distill-Qwen-1.5B model toward more concise reasoning, outperforming OpenAI’s o1-preview OpenAI (2024). By observing the training logs [^2] https://github.com/agentica-project/rllm of DeepScaleR in Figure 2, we find two issues:
@@
To this end, in this paper, we investigate how the model’s context length and the complexity of the training dataset influence the training process of R1-like reasoning models. Motivated by our observations, we propose FastCuRL, a simple yet efficient Curriculum Reinforcement Learning framework with a stage-wise context scaling strategy to improve the RL training efficiency and achieve concise CoT. Specifically, the proposed method alternates between CoT compression (long-to-short) and extension (short-to-long). We first compress CoT reasoning outputs, then extend them, and repeat this compress–extend cycle. The main goal of our method is to enhance CoT quality progressively. We hypothesize that this dynamic interplay fosters higher-quality reasoning by iteratively refining essential logical dependencies while pruning redundant or spurious steps. Meanwhile, during the iterative process, we control the complexity of the training samples, thereby ensuring training efficiency. Extensive experimental results demonstrate that our model FastCuRL-1.5B-V3 outperforms recent state-of-the-art reasoning baselines across five competition-level benchmarks, AIME 2024, AMC 2023, MATH 500, Minerva Math, and OlympiadBench. Furthermore, our model FastCuRL-1.5B-Preview surpasses DeepScaleR-1.5B-Preview on five competition-level benchmarks and only uses 50% training steps on a single node with 8 GPUs. We hope the findings presented in this paper, the models we have released, and the open-sourced code will benefit future research.
@@ 2 Methodology
Our method for achieving efficient LLM reasoning integrates three main components: (1) a resource-efficient RL algorithm, (2) a complexity-aware, mathematics-focused dataset, and (3) a stage-wise context scaling approach. This section provides a detailed description of the first two components.
@@ 2.1 Reinforcement Learning Algorithm
where $\beta$ denotes the coefficient of the KL penalty and the advantage $\mathrm{A}_{i}$ is computed from a group of rewards $\{r_{1},r_{2},...,r_{\mathrm{G}}\}$:
@@
To enhance the exploration capabilities of the policy model, we introduce an entropy bonus term,
@@
where $\mathbb{H}(\pi_{\theta}(o_{i}|q))$ indicates the entropy term and $\alpha$ denotes the entropy coefficient.
@@
2.2 Complexity-Aware Data Selection
@@ 3 Experiments
We designed a set of experiments, constrained by available computational resources, to investigate our central research question (Section 1): how do model context length and training data complexity affect the RL training of R1-like reasoning models? By analyzing the training behavior of small LLMs, we aim to extract practical insights. The objective is to move beyond simply providing empirical evidence of performance gains and to offer clear, actionable guidance for both academic and industrial implementation.
@@ 3.1 Experimental Setup
Table 3: Pass@1 performance across competition-level mathematical benchmarks.
3.2 Baselines
3.3 Benchmarks
3.4 Evaluation Metric
$\textsc{Pass@1}=\frac{1}{k}\sum_{i=1}^{k}p_{i},$ (4)
where $p_{i}$ is the correctness of the $i$-th response.
3.5 Main Processes and Results
In this section, we first validate the effectiveness of the complexity-aware data selection strategy. Then, we design a series of progressive experiments with varying context lengths and data complexities and analyze the experimental results.
@@ 3.5 Main Processes and Results
Figure 6: Entropy values of our models across all RL training stages. Here, data points are smoothed using a running average (window size = 10) and plotted at uniform intervals.
3.5.1 Main Results
Table 3 presents a comprehensive comparative analysis of the Pass@1 performance of our models against several state-of-the-art open-source baselines. The results unequivocally demonstrate the superiority of our proposed approach. Specifically, our final model, FastCuRL-1.5B-V3, achieves an average score of 61.6, significantly outperforming all baselines, including those with larger parameters. More specifically, FastCuRL-1.5B-V3 establishes new state-of-the-art results across all five competition-level benchmarks, scoring 90.5 on MATH 500, 49.6 on AIME 2024, 78.5 on AMC 2023, 34.7 on Minerva Math, and 54.5 on OlympiadBench. The consistent performance progression from FastCuRL-1.5B-Preview (57.5) to V3 (61.6) further validates the effectiveness and scalability of our training method.
In addition to overall performance, our models exhibit enhanced generalization capabilities. The initial FastCuRL-1.5B-Preview model already shows improvement over the DeepScaleR-1.5B-Preview on the AMC 2023 (74.2 vs. 73.6) and Minerva Math (31.6 vs. 30.2) test sets. This advantage is substantially amplified in our final model, FastCuRL-1.5B-V3, which widens the performance gap with scores of 78.5 and 34.7, respectively. From the perspective of training efficiency, these superior results are achieved with remarkable resource savings. Compared to the training regimen of the DeepScaleR-1.5B-Preview, our method utilizes only 50% of the training steps. It requires just a single node with 8 GPUs, thereby reducing training costs by more than half, as shown in Table 4. This demonstrates that our approach not only yields a more powerful model but also represents a more computationally efficient for developing advanced reasoning capabilities.
3.5.2 Analyzing Entropy in RL Training
Figure 6 presents the policy entropy and model performance across all training steps, segmented into five distinct stages (S-1 through S-5). The entropy trajectory exhibits three characteristic regimes: (i) gradual increase from 0.17 to 0.28 during initial exploration (S-1 to S-3), (ii) peak exploration at 0.40 during S-4, and (iii) substantial collapse to 0.22 in the final stage (S-5).
A central challenge in RL scaling is managing the exploration-exploitation trade-off, where a standard failure mode is "entropy collapse." This phenomenon occurs when the policy rapidly becomes deterministic, leading to a sharp decrease in entropy and causing the policy to prematurely converge to a sub-optimal local optimum, thereby losing its ability to discover more effective strategies. However, as depicted in Figure 6, our approach deviates significantly from this standard pattern. Instead of a premature collapse, the policy entropy demonstrates a prolonged and even increasing stage of exploration (S-2 to S-4), indicating that the policy actively maintains stochasticity. This sustained exploration is critical for navigating complex solution spaces and avoiding early stagnation.
By adjusting the reasoning context length and training data complexity, we effectively reinvigorate the policy’s exploratory drive at key training intervals. The pronounced entropy peak in S-4 exemplifies this effect, where the agent is stimulated to explore novel behaviors rather than settling for its current policy. Crucially, this period of heightened exploration lays the groundwork for a more effective exploitation stage in S-5. As the policy finally converges and entropy decreases, the model’s performance, measured by AIME 2024 Accuracy, continues to ascend to its peak. This demonstrates that our method successfully mitigates premature entropy collapse, enabling the model to achieve a significantly higher performance ceiling.
@@ 3.5.2 Analyzing Entropy in RL Training
Table 5: Statistics of the outputs of DeepSeek-R1-Distill-Qwen-1.5B and FastCuRL-1.5B-Preview.
@@
3.5.3 Dataset Complexity Verification
To validate the effectiveness of our dataset selection approach, we train three reasoning models under identical settings on L1, L2, and L3 with an 8K context length. As illustrated in Figure 5, the experimental results demonstrate consistent patterns across all three metrics that align with our expectations. The output length progressively increases from L1 to L3 throughout the training process, with models trained on more complex datasets generating substantially longer responses, reflecting the additional reasoning steps required for solving complex problems. Correspondingly, the clip ratio exhibits the exact ordering (L3 > L2 > L1), indicating that complex problems naturally demand longer solution outputs that more frequently exceed the context length. While the reward scores show an inverse relationship, with L1 achieving the highest scores due to its simpler problems, this pattern is expected. This observation demonstrates the inherent trade-off between problem complexity and output length, which supports our hypothesis that problem complexity directly correlates with required output length; more complex problems necessitate longer, more detailed reasoning outputs to arrive at correct solutions, thereby validating the importance of complexity-aware data selection.
3.5.4 Multi-Stage Experimental Results
In our experiments, we conduct three sets of multi-stage experiments, comprising 3, 4, and 5 training stages, respectively. The specific parameter settings are detailed in Table 3, and the corresponding results are presented in Table 2. In the first set of experiments, Exp-3 outperforms Exp-1 but requires additional training steps, as it is trained twice on the $\mathbf{L}$2 dataset. Considering the trade-off between performance and computational cost, we select Exp-1 as the base model for the second stage. Following the first stage, we observe that the average response length is between 6,000 and 7,000 tokens. This observation motivates our subsequent evaluation of various context lengths. As shown in Table 2, a 16K context length (Exp-6) yields the best performance, surpassing the longer 24K and 32K settings. Consequently, we select Exp-6 as the base model for the third stage.
Informed by the second set of experiments, we conduct a third set evaluating context lengths of 24K, 16K, and 8K. As shown in Table 2, the 16K context still achieves the best performance, though it shows negligible improvement over the fourth stage. Our analysis of this phenomenon suggests that during compression and extension, the model’s output length is initially constrained by the shorter context of the first stage. This constraint appears to compress the length of the thought process while enhancing its quality.
As the context length extensions in subsequent stages, the model begins to explore problems requiring longer CoT outputs. However, this expansion improves the model’s performance but also introduces repetitive thinking patterns. These patterns may not improve reasoning capabilities; on the contrary, they may diminish the model’s exploratory efficiency, particularly at excessively long context lengths. Therefore, we posit that subsequent context compression, as performed in the fourth stage, is necessary to improve CoT quality and enhance exploratory efficiency.
In the third set of experiments, we observe that neither increasing nor decreasing the context length from 16K yields comparable performance, suggesting that 16K represents an optimal configuration. This finding prompts a critical question: Does an optimal context length, or "sweet spot", exist for R1-like models? Specifically, we investigate whether 16K is the precise optimum for the DeepSeek-R1-Distill-Qwen-1.5B model or simply the best-performing value among the tested configurations of 8K and 24K.
To investigate this question, we conduct a comprehensive series of experiments examining model behavior across a finer-grained spectrum of context lengths. We train the model with various context length configurations while holding the entropy coefficient constant at $1\times 10^{-6}$ to observe entropy dynamics. As illustrated in Figure 8, our findings reveal a notable dichotomy. For context lengths of 4K, 8K, and 12K, the entropy exhibits a rapid decline, indicating a significant degradation in the model’s exploratory capabilities. In contrast, for context lengths of 16K, 20K, and 24K, the entropy demonstrates remarkable stability, converging to and maintaining a consistent value.
@@ 3.5.4 Multi-Stage Experimental Results
Figure 9: Performance comparison of training with and without Entropy bonus at 8k context length.
3.5.5 Analyzing Generated CoT Outputs
@@
Figure 7 compares FastCuRL-1.5B-Preview and DeepSeek-R1-Distill-Qwen-1.5B on the AIME 2024, measuring the average response length between correct and incorrect answers at the problem level to observe and analyse whether the long incorrect response is related to the difficulty of the problem. Across both models, incorrect answers (red bars) almost universally have greater average response lengths than correct answers (green bars). This suggests that models tend to generate more verbose content when producing incorrect answers, potentially reflecting "over-explanation" or "verbose reasoning" when the model is uncertain.
3.5.6 Ablation Study
Figure 9 presents a comprehensive comparison of model performance under different regularization configurations at an 8k context length. The upper panel demonstrates that models incorporating Entropy regularization (with Entropy & KL) or KL divergence alone (without Entropy) maintain consistently higher output lengths around 3,700-3,900 tokens throughout training, while the model without KL divergence regularization exhibits notably shorter outputs, stabilizing at approximately 3,000-3,500 tokens. This suggests that KL divergence plays a critical role in preventing output length collapse during long-context generation. The lower panel reveals that all three configurations achieve comparable reward scores, converging to approximately 0.55 after initial training steps, indicating that the choice of regularization strategy primarily affects generation length rather than quality. Notably, the combination of entropy bonus and KL divergence (orange line) demonstrates the most stable training dynamics across both metrics, suggesting potential synergistic effects between these two regularization techniques in balancing exploration and exploitation during policy optimization.
@@ 4 Related Work
Advancements in RL methodologies have considerably enhanced the reasoning capabilities of LLMs. A pivotal development in this domain is OpenAI’s o1 OpenAI (2024), which employs RL training to promote the development of long CoT reasoning in LLMs. This approach has significantly enhanced performance on complex mathematical and programming benchmarks. Building upon this foundation, DeepSeek-R1 DeepSeek-AI (2025) demonstrates that pure RL post-training via Group Reinforcement Policy Optimization (GRPO), without needing supervised pre-training, can directly perform robust CoT reasoning capabilities. Notably, this method not only achieves performance competitive with o1 but also exhibits emergent behaviors such as self-verification and multi-step planning. Building on these advancements, the research community has been collectively working to study and apply the methodology of DeepSeek-R1 to enhance the CoT reasoning capabilities of various sizes of language models, yielding remarkable progress, such as Face (2025); Luo et al. (2025); Zeng et al. (2025); Liu et al. (2025); Yu et al. (2025); Song and Zheng (2025). Different from the existing approaches, we employ an iterative stage-wise scaling method that alternates between CoT compression (long-to-short) and CoT extension (short-to-long) training stages. We initially compress the CoT outputs, subsequently perform extension, followed by re-compression and re-extension in a cyclical manner. Specifically, we hypothesize that this dynamic interplay between compression and extension facilitates the emergence of higher-quality reasoning responses by iteratively refining the essential logical dependencies while eliminating redundant or spurious CoT reasoning outputs.